Assume CGM Corp has issued and outstanding 22,000 shares of 5%, $25 par value preferred stock and issued and outstanding 15,000 shares of $15 stated value common stock. The board of directors declares dividends of $50,000, the preferred stock is cumulative, and $17,500 dividends is in arrears. What amount is allocated to common stockholders?
Answer #1

Current year preferred dividend = 22000*25*5% = 27500

Amount allocated to preferred stockholders = Dividend in arrears+Current year dividend

= 17500+27500

Amount allocated to preferred stockholders = 45000

Amount allocated to common Stockholders = 50000-45000 = $5000